What is the price of a STRIPS maturing in 18 years with a face value of $15,000 and yield to maturity of 6.50%. The current rate on S-T Federal T-Bills is 4.75%. YOU MUST SHOW ALL WORK TO RECEIVE CREDIT. MAKE SURE YOU USE AT LEAST 4 DECIMAL PLACES IN ALL CALCULATIONS.
Answer #1

Bond price=(Cash flow in year 1)/(1+YTM)^1+(Cash flow in year 2)/(1+YTM)^2+(Cash flow in year 3)/(1+YTM)^3+(Cash flow in year 4)/(1+YTM)^4+(Cash flow in year 5)/(1+YTM)^5+(Cash flow in year 6)/(1+YTM)^6+(Cash flow in year 7)/(1+YTM)^7+(Cash flow in year 8)/(1+YTM)^8+(Cash flow in year 9)/(1+YTM)^9+(Cash flow in year 10)/(1+YTM)^10+(Cash flow in year 11)/(1+YTM)^11+(Cash flow in year 12)/(1+YTM)^12+(Cash flow in year 13)/(1+YTM)^13+(Cash flow in year 14)/(1+YTM)^14+(Cash flow in year 15)/(1+YTM)^15+(Cash flow in year 16)/(1+YTM)^16+(Cash flow in year 17)/(1+YTM)^17+(Cash flow in year 18)/(1+YTM)^18+ Face value/(1 + YTM)^18

The cash flows are the coupon payments and here the coupon payment = (Coupon rate)*(Face value)=0%*15000=0
So, the bond price=Face value/(1 + YTM)^18=15000/(1+6.5%)^18=$4828.3453